# Motion of spinning particles and chaos bound in Reissner-Nordström spacetime

### v1

@@ 3.3 Charged particle
When the particle carries an electric charge, the electromagnetic force exerted on it by the BH affects its position of the unstable equilibrium orbit as well as the value of the exponent. In this regard, we carry out an investigation into the influences exerted by the particle’s charge, spin and total angular momentum on the exponent. We also use Eqs. (9), (10), (25), (26), (27) and (35) to calculate the exponent and generate corresponding plots. Based on the computation presented in the preceding subsection, we continue to set $Q=0.95$ in this analysis.
@@
We initially examine, through the variation of the particle’s charge, whether the chaos bound is violated in Figure 5. As shown in the figure, when the charge remains below a specific critical threshold, the exponent consistently exceeds the surface gravity, thereby violating the bound. When the spin and angular momentum direction are aligned, it can be observed that, for a given charge, a larger spin magnitude corresponds to a higher value of the associated exponent. This situation also applies when the two directions are opposite. This implies that an increase in spin magnitude intensifies the chaotic behavior of the particle. Furthermore, it is shown here that both scenarios, where the charge-to-mass ratio of the particle is greater than $1.00$ or less than $1.00$, can lead to the bound’s violations. Given this, we order $q=0.70$ in the subsequent calculations.
@@
The relationship between the effective potential and radius of the charged particle is plotted in Figure 6. This figure bears similarities to Figure 2, namely, the maxima of the potential all exhibit an increasing trend as the spin rises. Nevertheless, owing to the comparatively weak electromagnetic force, the discrepancy in this particular scenario isn’t large. For a given spin, these maxima of the potential for the charged particle are larger than those for the neutral particle. Consequently, for the particle with the same spin, the unstable equilibrium orbit of the charged particle is farther from the event horizon compared to that of the uncharged particle.
@@
The variation of the exponent with respect to the spin is depicted in Figure 7. In this figure, two situations still persist: one in which the spin direction is aligned with the direction of the angular momentum, and the other in which they are anti-aligned. When the two directions are aligned, the trend of the exponent changing with the spin is similar to that in Figure 3. However, the threshold values of the angular momentum corresponding to the violation of the bound are not identical. For $L=15.00$, the bound for the spinless particle is not violated. When the two directions are anti-aligned, except for the case of $L=5.00$, the variation of the exponent with respect to the spin magnitude is similar to that shown in Figure 3. We first elaborate on the situation associated with this angular momentum. In this scenario, the exponent has three minima. As the spin magnitude increases, it first decreases and then increases, presenting a minimum value smaller than the surface gravity when S=0.50S=-0.50. The spin magnitude corresponding to the maximum value of the exponent has changed, showing a decrease compared to the spin magnitude in Figure 3. The possible cause of the aforementioned phenomenon might be that the electromagnetic force is not sufficiently strong. Consequently, when the electromagnetic force is relatively weak and the angular momentum is greater than $5.00$, this electromagnetic force cannot alter the trend of how the exponent varies with respect to the spin. Instead, it can only affect the magnitude of the exponent, which, in turn, influences the spin’s threshold for violating the bound.
@@
Figure 8 illustrates the variation of the exponent with respect to the total angular momentum of the charged particle. This analysis is carried out under two separate scenarios: when the spin is aligned with the angular momentum direction and when the two directions are opposite. In the scenario of alignment, for a given fixed spin magnitude, the exponent increases as the angular momentum rises. However, the angular momentum must surpass a specific threshold for the exponent’s value to exceed the surface gravity, thereby resulting in a violation of the bound. Consequently, an increase in the angular momentum intensifies the chaotic behavior of the particle. The values of the exponent under specific angular momentum and spin conditions are generally smaller than those depicted in Figure 4 for the corresponding cases. When the two directions are opposite, for a fixed spin magnitude, the exponent still exhibits an increasing trend with the rise in the angular momentum. Nevertheless, at the spin magnitude of $2.00$, a violation of the bound occurs irrespective of the angular momentum values. For other spin magnitudes, violations only manifest when the angular momentum exceeds certain thresholds.
@@ 4 Conclusions and discussions
In this work, we investigated the motion of the spin particles around the RN BH, and discussed the relationship between the LE of its classical chaotic motion and the chaos bound. Taking into account the two cases where the spin direction of the particle is aligned and anti-aligned with its total angular momentum direction, we found that the exponent can surpass the bound under the specific conditions. Therefore, the phenomenon of violations of the bound is also observable within the spinor field.
For the neutral particle, we initially fixed its total angular momentum and analyzed the influence of the BH’s charge on the bound. It was discovered that the violations of the bound occurs only when the charge exceeds its threshold values, and the different spins correspond to the distinct charge thresholds. Based on these thresholds, we further fixed the charge and examine the impacts of the particle’s spin and angular momentum on the exponent. For a given angular momentum, when the spin aligns with the angular momentum direction, the exponent increases with the rising spin. Conversely, when their directions are opposite, the exponent displays two maxima as the spin magnitude increases.
When the particle is charged, we took into account the influence of the electromagnetic force. It was observed that there are certain similarities in the variation of the exponent with the spin between the charged and neutral particles (except for the case where the angular momentum is $5.00$). This indicates that the electromagnetic force cannot alter the trend of how the exponent varies with respect to the spin or angular momentum, but it can modify the exponent’s value. The likely reason for this phenomenon is that the electromagnetic force is not sufficiently strong. If both the BH’s and particle’s charges are large enough simultaneously, the trend of the exponent’s variation may change.
There are two explanations for the violation of the bound. The first explanation asserts that there is no real violation of the bound. Instead, it requires modification HT5 ; LTW1 . The second explanation admits the violation of the bound, and its proponents posit that this violation may be associated with the thermodynamic stability of BHs. However, the subsequent research has demonstrated that when the BH undergoes the phase transition, the violation of the bound occurs in the BH branch within its stable region LGD .

### v2

@@ 3.3 Charged particle
When the particle carries an electric charge, the electromagnetic force exerted on it by the BH affects its position of the unstable equilibrium orbit as well as the value of the exponent. In this regard, we carry out an investigation into the influences exerted by the charge, spin and total angular momentum of the particle on the exponent. We also use Eqs. (9), (10), (25), (26), (27) and (35) to calculate the exponent and generate plots. Based on the computation presented in the preceding subsection, we continue to set $Q=0.95$ in this analysis.
@@
We initially examine, through the variation of the particle’s charge, whether the bound is violated in Figure 5. As shown in the figure, when the charge remains below a specific critical threshold, the exponent consistently exceeds the surface gravity, thereby violating the bound. When the spin and angular momentum directions are aligned, it can be observed that, for a given charge, a larger spin magnitude corresponds to a higher value of the exponent. This situation also applies when the two directions are opposite. This implies that an increase in the spin magnitude intensifies the chaotic behavior of the particle. Furthermore, it is shown here that both scenarios, where the charge-to-mass ratio of the particle is greater than $1.00$ or less than $1.00$, can lead to violations of the bound. Given this, we order $q=0.70$ in the subsequent calculations.
@@
The relationship between the effective potential and radius of the charged particle is plotted in Figure 6. This figure bears similarities to Figure 2, namely, the maxima of the potential all exhibit an increasing trend as the spin rises. Nevertheless, owing to the comparatively weak electromagnetic force, the discrepancy in this particular scenario isn’t large. For a given spin, these maxima of the potential for the charged particle are larger than those for the neutral particle. Consequently, for the particle with the same spin, the unstable equilibrium orbit of the charged particle is farther from the event horizon compared to that of the neutral particle.
@@
The variation of the exponent with respect to the spin is depicted in Figure 7. In this figure, two situations still persist: one in which the spin direction is aligned with the angular momentum direction, and the other in which they are anti-aligned. When the two directions are aligned, the trend of the exponent changing with the spin is similar to that in Figure 3. However, the threshold values of the angular momentum corresponding to the violation of the bound are not identical. For $L=15.00$, the bound for the spinless particle is not violated. When the two directions are anti-aligned, except for the case of $L=5.00$, the variation of the exponent with respect to the spin magnitude is similar to that shown in Figure 3. We first elaborate on the situation associated with this angular momentum. In this scenario, the exponent has three minima. As the spin magnitude increases, it first decreases and then increases, presenting a minimum value smaller than the surface gravity when S=0.50S=-0.50. The spin magnitude corresponding to the maximum value of the exponent has changed, showing a decrease compared to the spin magnitude in Figure 3. The possible cause of the aforementioned phenomenon might be that the electromagnetic force is not sufficiently strong. Consequently, when the electromagnetic force is relatively weak and the angular momentum is greater than $5.00$, this electromagnetic force cannot alter the trend of how the exponent varies with respect to the spin. Instead, it can only affect the magnitude of the exponent, which, in turn, influences the spin’s threshold for violating the bound.
@@
Figure 8 illustrates the variation of the exponent with respect to the total angular momentum of the charged particle. This analysis is carried out under two separate scenarios: when the spin direction is aligned with the angular momentum direction and when the two directions are opposite. In the scenario of alignment, for a given fixed spin magnitude, the exponent increases as the angular momentum rises. However, the angular momentum must surpass a specific threshold for the exponent’s value to exceed the surface gravity, thereby resulting in a violation of the bound. Consequently, an increase in the angular momentum intensifies the chaotic behavior of the particle. The values of the exponent under specific angular momentum and spin conditions are generally smaller than those depicted in Figure 4 for the corresponding cases. When the two directions are opposite, for a fixed spin magnitude, the exponent still exhibits an increasing trend with the rise in the angular momentum. Nevertheless, at the spin magnitude of $2.00$, a violation of the bound occurs irrespective of the angular momentum values. For other spin magnitudes, violations only manifest when the angular momentum exceeds certain thresholds.
@@ 4 Conclusions and discussions
In this work, we investigated the motion of the spinning particles around the RN BH, and discussed the relationship between the LEs of their classical chaotic motion and the chaos bound. Taking into account the two cases where the spin direction of the particle is aligned and anti-aligned with its total angular momentum direction, we found that the exponents can surpass the bound under the specific conditions. Therefore, the phenomenon of the chaos bound violation is also observable within the spinor field.
For the neutral particle, we initially fixed its total angular momentum and analyzed the influence of the BH’s charge on the bound. It was discovered that the violation occurs only when the charge exceeds its threshold values, and the different spins correspond to the distinct charge thresholds. Based on these thresholds, we further fixed the charge and examine the impacts of the spin and angular momentum on the exponent. For a given angular momentum, when the spin direction aligns with the angular momentum direction, the exponent increases with the rising spin. Conversely, when their directions are opposite, the exponent displays two maxima as the spin magnitude increases.
When the particle is charged, we took into account the influence of the electromagnetic force. It was observed that there are certain similarities in the variation of the exponent with the spin between the charged and neutral particles (except for the case where the angular momentum is $5.00$). This indicates that the electromagnetic force cannot alter the trend of how the exponent varies with respect to the spin and angular momentum, but it can modify the value of the exponent. The likely reason for this phenomenon is that the electromagnetic force is not sufficiently strong. If both the BH’s and particle’s charges are large enough simultaneously, the trend of the exponent’s variation may change.
There are two explanations for the violation of the bound. The first explanation asserts that there is no real violation. Instead, it requires a modification HT5 ; LTW1 . The second explanation admits the violation, and its proponents posit that this violation may be associated with the thermodynamic stability of BHs. However, the subsequent research has demonstrated that when the BH undergoes the phase transition, the violation occurs in the BH branch within its stable region LGD .